divide 184 in two parts as one third of one part may exceed to one seventh to other by 7. Find the two parts?
Answer by 03008082502(3)

1st part =x
2nd part =184-x
according to the condition
x /3=(184-x)/7 +7
after slove this equation we get
x =114.9